Part A: ΔH∘rxn= 121 kJ ; ΔS∘rxn=− 246 J/K ; T= 291 K Express your answer...

Part A:

ΔH∘rxn= 121 kJ ; ΔS∘rxn=− 246 J/K ; T= 291 K

Express your answer as an integer.

Part B:

ΔH∘rxn=− 121 kJ ; ΔS∘rxn= 246 J/K ; T= 291 K

Express your answer as an integer.

Part C:

ΔH∘rxn=− 121 kJ ; ΔS∘rxn=− 246 J/K ; T= 291 K

Express your answer as an integer.

Part D:

ΔH∘rxn=− 121 kJ ; ΔS∘rxn=− 246 J/K ; T= 600 K

Express your answer as an integer.

Part E:

Predict whether or not the reaction in part A will be spontaneous at the temperature indicated.

Part F:

Predict whether or not the reaction in part B will be spontaneous at the temperature indicated.

Part G:

Predict whether or not the reaction in part C will be spontaneous at the temperature indicated.

Part H:

Predict whether or not the reaction in part D will be spontaneous at the temperature indicated.
Answer #1

A)

ΔHo = 121 KJ

ΔSo = -246 J/K

= -0.246 KJ/K

T = 291 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = 121.0 - 291.0 * -0.246

ΔGo = 192.586 KJ

Answer: 193 KJ

B)

ΔHo = -121 KJ

ΔSo = 246 J/K

= 0.246 KJ/K

T = 291 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -121.0 - 291.0 * 0.246

ΔGo = -192.586 KJ

Answer: -193 KJ

C)

ΔHo = -121 KJ

ΔSo = -246 J/K

= -0.246 KJ/K

T = 291 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -121.0 - 291.0 * -0.246

ΔGo = -49.414 KJ

Answer: -49 KJ

D)

ΔHo = -121 KJ

ΔSo = -246 J/K

= -0.246 KJ/K

T = 600 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -121.0 - 600.0 * -0.246

ΔGo = 26.6 KJ

Answer: 27 KJ

E)

Since ΔGo is positive, the reaction would be non spontaneous

Answer: non spontaneous

F)

Since ΔGo is negative, the reaction would be spontaneous

Answer: spontaneous

G)

Since ΔGo is negative, the reaction would be spontaneous

Answer: spontaneous

H)

Since ΔGo is positive, the reaction would be non spontaneous

Answer: non spontaneous